# Exact path integrals on half-line in quantum cosmology with a fluid clock and aspects of operator ordering ambiguity

### v1

@@ V.1 Case I: Flat FLRW spacetime with a perfect fluid as clock
$\int_{0}^{\infty}{\rm d}s_{\sf f}\,{\rm e}^{\frac{i}{\hbar}\lambda s_{\sf f}}=\pi\delta(\lambda/\hbar)+\mathcal{P}\left(\frac{i}{\lambda/\hbar}\right).$ (99)
Following this we can immediately perform the integration over $\lambda$, that is,
@@ V.2 Case II: Bianchi I spacetime with a free massless scalar field and a perfect fluid as clock
We have seen previously that the quantization of one-dimensional minisuperspace does not respect the conformal invariance present in the classical theory. Now we should look for a higher dimensional example of minisuperspace. The simplest extension from the above example would be to include a free massless scalar field, and the corresponding minisuperspace $(a,\phi)$ will have $D=2$. Even though this is a valid example, notice that in this case the conformal term $+\hbar^{2}\frac{D-2}{8(D-1)}\mathcal{R}$ vanishes identically and no additional term is required for conformal invariance. For this reason, we shall give an example of another even higher dimensional minisuperspace so that the role and relevance of this potential-class term become clear. We shall, in particular, consider the Bianchi I spacetime, which describes a spatially flat, homogeneous, but anisotropic universe. Notice that the action in 74, can be obtained from by starting with
@@
where we have introduced $\pm$ sign in the fixed momentum state for anistorpy and scalar field reflecting the fact that integral over $\bm{\Pi}$ and $\mathsf{p}_{\phi}$ in V.2 is invariant under the transformations $\bm{\Pi}\to-\bm{\Pi}$ and/or $\mathsf{p}_{\phi}\to-\mathsf{p}_{\phi}$. We also note that the above wave function has the correct normalization with respect to the inner product 78. Furthermore, this wave function, obtained from the path integral, satisfies the Wheeler-DeWitt equation 77. Therefore, the path integral and canonical approach produce consistent results. As expected, the wave function follows the scaling property under lapse rescaling 50, and the system is invariant under such rescalings as discussed for the canonical case in Sec. IV.2. Notice the fact that the quantum correction to the potential ($+\frac{D-2}{8(D-1)}\mathcal{R}$) that appeared in the Wheeler-DeWitt equation is crucial to this ambiguity-free description. It can be checked that without this additional correction, the index of the Bessel function depends on the ambiguity parameter $p$, making the predictions of the theory essentially dependent on the operator ordering through the parameter $p$.
Thus, the invariance of the systems with minisuperspace of dimension $D\geq 2$ under lapse rescaling implies the operator ordering ambiguities (conformal and factor classes) leave no physical imprint for these systems. Notice that for this conclusion to hold the potential-class ambiguity term has to be fixed and is not arbitrary. Therefore, one may conclude that building quantum theories which preserve the symmetries of the classical theory, under point canonical transformation and lapse rescalings, has resulted into the ambiguities being immaterial for the physical predictions.
@@ VI Discussion and Conclusion
We see that the ambiguities in the Wheeler-DeWitt equation can be characterized by three distinct functions, which we have called conformal-class, factor-class, and potential-class ambiguities. Among these three classes, the factor-class only results in an overall factor in the wave function and gets canceled in the inner products, leaving no imprint in the physical predictions. Whereas, the other two ambiguities require careful treatment. We further find that if we demand symmetry under point canonical transformations or the coordinate transformations in the minisuperspace, then the Laplacian in the Wheeler-DeWitt equation is defined with the Laplace-Beltrami operator in curved spaces depending only on the metric of the minisuperspace. However, this does not fix any ambiguity in the Hamiltonian as we still have the lapse rescaling freedom in the classical theory. We have shown that two equivalent classical theories which differ by lapse rescaling lead to two distinct and inequivalent quantum theories which differ by operator ordering ambiguity of the conformal-class. Therefore, resolving conformal-class ambiguity is equivalent to demanding conformal invariance in the quantum theory. This is achieved by choosing a particular form of the potential-class ambiguity term proportional to the Ricci scalar of the minisuperspace, more specifically $+\frac{D-2}{8(D-1)}\mathcal{R}$. Once this choice has been made, surprisingly, all the ambiguities of the conformal-class become irrelevant as these no longer leave any imprint on the inner products between quantum states. However, this symmetry principle does not work for $D=1$ case and only works for the cases $D\geq 2$.
@@
We have derived the above results in both the canonical and path integral approaches. Noticing that operator ordering ambiguity and the problem of time in the path integral approach are rarely discussed in the literature, in this work, we have computed exact path integrals for a flat and homogeneous quantum universe, both having isotropic FLRW metric and anisotropic Bianchi I metric and containing a single fluid with an arbitrary equation of state parameter $\omega$. The fluid acts as a reference clock and enables the unitary evolution of the quantum state of the universe. In the case of anisotropic cosmology, we have shown that a massless scalar field can easily be incorporated as well, and still, an exact path integral can be performed. As the path integrals are not of the Gaussian or canonical form, we had to introduce coordinate transformations to bring them into forms, for which the evaluation of the path integral is known. All the coordinate transformations required a careful treatment of the Jacobian factors in the integral and quantum corrections to the potential. We have further verified that the wave functions obtained from the path integral formalism, indeed, satisfy the corresponding Wheeler-DeWitt equations or the quantum Hamiltonian constraint. Moreover, the path integral computations produce wave functions with correct normalization. These concrete examples support our general derivations of the results stated above.
Further, we note that in a timeless formalism of quantum cosmology, the notion of unitary quantum evolution and inner products is not well defined. As a result, even though Halliwell noticed that the potential-class term could be wisely chosen to resolve the conformal-class ambiguities [7], the author could not discuss the inner product between the quantum states, which we have done here. Moreover, our work is different in at least another key aspect: we have extended the proof of the lapse choice invariance to the case of path integral formalism, which is absent in [7].
@@
The ordering ambiguity still remains in the case of $D=1$ minisuperspace, and it cannot be removed, at least with the techniques we discussed in this article. More specifically, we have shown that the wave functions derived in both the isotropic and anisotropic cases have a Besselian form. In the case of the higher dimensional minisuperspace (in our example $D=4$), the index of the Bessel function is independent of the ambiguity parameter $p$, which only appears in the power of the scale factor multiplied with the Bessel function and does not appear in the inner products as this factor is canceled by appropriate integration measure. Therefore, the ambiguity is removed from the physical predictions. On the other hand, in the case of $D=1$ minisuperspace, the index of the Bessel function depends on the ambiguity parameter and cannot be removed by the integration measure. As a result, the physical predictions of the theory will always depend on the ordering choice of the Hamiltonian, as depicted in [17, 20].
@@ acknowledgements
Research of KL is partially supported by SERB, Govt. of India through a MATRICS research grant no. MTR/2022/000900. The research of VM is funded by the INSPIRE fellowship from the DST, Government of India (Reg. No. DST/INSPIRE/03/2019/001887). The research of HSS is supported by the Core Research Grant CRG/2021/003053 from the Science and Engineering Research Board, India. We are thankful to Sumanta Chakraborty for insightful discussions on the problem. VM acknowledges Kaustav Das for helpful discussions during the preparation of this manuscript.

### v2

@@ V.1 Case I: Flat FLRW spacetime with a perfect fluid as clock
$\int_{0}^{\infty}{\rm d}s_{\sf f}\,{\rm e}^{\frac{i}{\hbar}\lambda s_{\sf f}}=\pi\delta(\lambda/\hbar)+\mathcal{P}\left(\frac{i}{\lambda/\hbar}\right),$ (99)
where $\mathcal{P}$ denotes the Cauchy principal value. Following this we can immediately perform the integration over $\lambda$, that is,
@@ V.2 Case II: Bianchi I spacetime with a free massless scalar field and a perfect fluid as clock
We have seen previously that the quantization of one-dimensional minisuperspace does not respect the conformal invariance present in the classical theory. Now we should look for a higher dimensional example of minisuperspace. The simplest extension from the above example would be to include a free massless scalar field, and the corresponding minisuperspace $(a,\phi)$ will have $D=2$. Even though this is a valid example, notice that in this case the conformal term $+\hbar^{2}\frac{D-2}{8(D-1)}\mathcal{R}$ vanishes identically and no additional term is required for conformal invariance. For this reason, we shall give an example of another even higher dimensional minisuperspace so that the role and relevance of this potential-class term become clear. We shall, in particular, consider the Bianchi I spacetime, which describes a spatially flat, homogeneous, but anisotropic universe. Notice that the action in 74, can be obtained by starting with
@@
where we have introduced $\pm$ sign in the fixed momentum state for anisotropy and scalar field reflecting the fact that integral over $\bm{\Pi}$ and $\mathsf{p}_{\phi}$ in V.2 is invariant under the transformations $\bm{\Pi}\to-\bm{\Pi}$ and/or $\mathsf{p}_{\phi}\to-\mathsf{p}_{\phi}$. We also note that the above wave function has the correct normalization with respect to the inner product 78. Furthermore, this wave function, obtained from the path integral, satisfies the Wheeler-DeWitt equation 77. Therefore, the path integral and canonical approach produce consistent results. As expected, the wave function follows the scaling property under lapse rescaling 50, and the system is invariant under such rescalings as discussed for the canonical case in Sec. IV.2. Notice the fact that the quantum correction to the potential ($+\frac{D-2}{8(D-1)}\mathcal{R}$) that appeared in the Wheeler-DeWitt equation is crucial to this ambiguity-free description. It can be checked that without this additional correction, the index of the Bessel function depends on the ambiguity parameter $p$, making the predictions of the theory essentially dependent on the operator ordering through the parameter $p$.
Thus, the invariance of the systems with minisuperspace of dimension $D>2$ under lapse rescaling implies the operator ordering ambiguities (conformal and factor classes) leave no physical imprint for these systems. Notice that for this conclusion to hold the potential-class ambiguity term has to be fixed and is not arbitrary. Therefore, one may conclude that building quantum theories which preserve the symmetries of the classical theory, under point canonical transformation and lapse rescalings, has resulted into the ambiguities being immaterial for the physical predictions.
@@ VI Discussion and Conclusion
We see that the ambiguities in the Wheeler-DeWitt equation can be characterized by three distinct functions, which we have called conformal-class, factor-class, and potential-class ambiguities. Among these three classes, the factor-class only results in an overall factor in the wave function and gets canceled in the inner products, leaving no imprint in the physical predictions. Whereas the other two ambiguities require careful treatment. We further find that if we demand symmetry under point canonical transformations or the coordinate transformations in the minisuperspace, then the Laplacian in the Wheeler-DeWitt equation is defined with the Laplace-Beltrami operator in curved spaces, depending only on the metric of the minisuperspace. However, this does not fix any ambiguity in the Hamiltonian, as we still have the lapse rescaling freedom in the classical theory. We have shown that two equivalent classical theories, which differ by lapse rescaling, lead to two distinct and inequivalent quantum theories which differ by operator ordering ambiguity of the conformal-class. Therefore, resolving conformal-class ambiguity is equivalent to demanding conformal invariance in the quantum theory. This is achieved by choosing a particular form of the potential-class ambiguity term proportional to the Ricci scalar of the minisuperspace, more specifically $+\frac{D-2}{8(D-1)}\mathcal{R}$. Once this choice has been made, surprisingly, all the ambiguities of the conformal-class become irrelevant as these no longer leave any imprint on the inner products between quantum states. However, this symmetry principle does not work for $D=1,2$ case and only works for the cases $D>2$. Only a specific choice for the potential-class term results in a conformally invariant theory in which conformal-class ambiguities become irrelevant. However, even a small perturbation of the potential-class term destroys the conformal invariance of the quantum theory and therefore, the ambiguities remain at large. Therefore, even though it is reasonable to preserve the classical symmetries in the quantum theory, and thereby arrive at a unique prescription for writing the quantum Hamiltonian, the operator ordering ambiguity cannot be resolved fully without empirical input.
@@
We have derived the above results in both the canonical and path integral approaches. Noticing that operator ordering ambiguity and the problem of time in the path integral approach are rarely discussed in the literature, in this work, we have computed exact path integrals for a flat and homogeneous quantum universe, both having isotropic FLRW metric and anisotropic Bianchi I metric and containing a single fluid with an arbitrary equation of state parameter $\omega$. The fluid acts as a reference clock and enables the unitary evolution of the quantum state of the universe. In the case of anisotropic cosmology, we have shown that a massless scalar field can easily be incorporated as well, and still, an exact path integral can be performed. As the path integrals are not of the Gaussian or canonical form, we had to introduce coordinate transformations to bring them into forms, for which the evaluation of the path integral is known. All the coordinate transformations required a careful treatment of the Jacobian factors in the integral and quantum corrections to the potential. We have further verified that the wave functions obtained from the path integral formalism, indeed, satisfy the corresponding Wheeler-DeWitt equations or the quantum Hamiltonian constraint. Moreover, the path integral computations produce wave functions with correct normalization. These concrete examples support our general derivations of the results stated above. Even though Halliwell noticed that the potential-class term could be wisely chosen to resolve the conformal-class ambiguities [7], the author did not discuss the inner product between the quantum states, which we have done here. This question can be addressed in the timeless formalism of quantum cosmology considered in that analysis [7] by introducing the notion of Hilbert space and inner product by group averaging method [97]. Moreover, our work is different in at least another key aspect: we have extended the proof of the lapse choice invariance to the case of the path integral formalism, which is absent in [7].
@@
The ordering ambiguity still remains in the case of $D=1$ and $D=2$ minisuperspace, and it cannot be removed, at least with the techniques we discussed in this article. More specifically, we have shown that the wave functions derived in both the isotropic and anisotropic cases have a Besselian form. In the case of the higher dimensional minisuperspace (in our example $D=4$), the index of the Bessel function is independent of the ambiguity parameter $p$, which only appears in the power of the scale factor multiplied with the Bessel function and does not appear in the inner products as this factor is canceled by appropriate integration measure. Therefore, the ambiguity is removed from the physical predictions. On the other hand, in the case of $D=1$ minisuperspace, the index of the Bessel function depends on the ambiguity parameter and cannot be removed by the integration measure. As a result, the physical predictions of the theory will always depend on the ordering choice of the Hamiltonian, as depicted in [17, 20].
@@ acknowledgements
Research of KL is partially supported by SERB, Govt. of India through a MATRICS research grant no. MTR/2022/000900. The research of VM is funded by the INSPIRE fellowship from the DST, Government of India (Reg. No. DST/INSPIRE/03/2019/001887). The research of HSS is supported by the Core Research Grant CRG/2021/003053 from the Science and Engineering Research Board, India. We are thankful to Sumanta Chakraborty for insightful discussions on the problem. VM acknowledges Kaustav Das for helpful discussions during the preparation of this manuscript. HSS and VM are grateful to the organizers of the workshop IAGRG School on Gravitation and Cosmology (code: ICTS/IAGRG2023/10) and the International Centre for Theoretical Sciences (ICTS) for hospitality.